OpenStudy (anonymous):

find the distance of the line 5x-12y = 26 from (1,5)

jagr2713 (jagr2713):

Hi there, here is the well known formula for the distance of a point (x,y) from a line: \[Distance(ax+by+c=0,(x_{0}y _{0})):\] So in our case the point is (1,5) and the line is : 5x-12y-26=0 so directly putting values in the formula we get: \[distance=\frac{ \left| 5x1-12x5-26 \right| }{ \sqrt{(5^{2}+12^{2})} }\]

jagr2713 (jagr2713):

Can you solve from here @martaamador62

jagr2713 (jagr2713):

Seems like you left but i will finish it off for you :D So we solve \[Distance=\frac{ \left| 5-60-26 \right| }{ \sqrt{(169)} }\] 5-60=-55 -55-26=-81 sqrt 169=13 \[distance=\frac{ \left| -81 \right| }{ 13}=\frac{ 81 }{ 13 }\]

OpenStudy (anonymous):

Thank you @jagr2713 :) and i'm so sorry i didn't immediately responded to this question i fell asleep :)))